\begin{thebibliography}{99}
%
%
%
%
\bibitem{Ar} M. Artin,
Supersingular $K3$ surfaces.
Ann. Sci. \'Ecole Norm. Sup. (4) {\bf 7} (1974), 543--567 (1975).
\bibitem{Artin1962}
M.~Artin,
Some numerical criteria for contractability of curves on algebraic surfaces,
Am.~J.~Math. {\bf 84} (1962), 485--496.
%
%
\bibitem{Artin1966}
M.~Artin,
On isolated rational singularities of surfaces,
Am.~J.~Math. {\bf 88} (1966), 129--136.
%
%
\bibitem{CS} J. H. Conway and N. J. Sloane,
Sphere packings, lattices and groups. Third edition. With additional contributions by E. Bannai, R. E.
Borcherds, J. Leech, S. P. Norton, A. M. Odlyzko, R. A. Parker, L. Queen and B. B. Venkov. Grundlehren der
Mathematischen Wissenschaften, {\bf 290}. Springer-Verlag, New York, 1999.
%
%
\bibitem{CD}
F.~Cossec and I.~Dolgachev,
Enriques surfaces. I.
Progress in Mathematics, 76.
Birkh\"auser Boston, Inc., Boston, MA, 1989.
%
%
\bibitem{DK} I. Dolgachev and J. Keum,
Finite groups of symplectic automorphisms of K3 surfaces in positive characteristic,
Ann. \ of Math. \ {\bf 169} (2009), 269?-313.
%math.AG/0403478.
%
%
\bibitem{DKondo}
I.~Dolgachev and S.~Kondo,
A supersingular $K3$ surface in characteristic $2$ and the Leech lattice,
Int. Math. Res. Not. 2003, no. 1, 1--23.
%
%
%
%
%
%
\bibitem{Goto}
Y.~Goto, \emph{On the {N}\'eron-{S}everi groups of some {$K3$} surfaces},
  The arithmetic and geometry of algebraic cycles (Banff, AB, 1998), CRM Proc.
  Lecture Notes, vol.~24, Amer. Math. Soc., Providence, RI, 2000, pp.~305--328.
%
%
\bibitem{Ko}
S.~Kondo,
Automorphisms of algebraic $K3$ surfaces which act trivially on Picard groups.
J. Math. Soc. Japan 44 (1992), no. 1, 75--98.
%
%
\bibitem{Nikulin} V. V. Nikulin,
Integer symmetric bilinear forms and some of their applications.
Math. USSR Izvestija {\bf 14} (1980), no.~1, 103--167.
%
%
\bibitem{Ni} V. V. Nikulin, Weil linear systems on singular $K3$ surfaces.
Algebraic geometry and analytic geometry (Tokyo, 1990), 138--164,
ICM-90 Satell. Conf. Proc.
%
%
\bibitem{Ogus}
A.~Ogus,
A crystalline Torelli theorem for supersingular $K3$ surfaces,
Arithmetic and geometry,
Vol II, pp.~361--349,
Progr.~Math.~36, Birkh\"auser, Boston, 1983.
%
%
\bibitem{RS1} A.~N.~Rudakov and I.~R.~Shafarevich,
Supersingular $K3$ surfaces over fields of characteristic $2$.
Math. USSR-Izv. {\bf 42} (1978), no. 4, 147--165 (1979).
%
%
\bibitem{RS2} A.~N.~Rudakov and I.~R.~Shafarevich,
Surfaces of type $K3$ over fields of finite characteristic. (Russian)
Current problems in mathematics, Vol. {\bf 18}, pp. 115--207,
Akad. Nauk SSSR, Vsesoyuz. Inst. Nauchn. i Tekhn. Informatsii, Moscow, 1981.
%
%
\bibitem{SD} B. Saint-Donat,
Projective models of $K-3$ surfaces.
Amer. J. Math. {\bf 96} (1974), 602--639.
%
%
%
\bibitem{ShimadaEll} I. Shimada,
On elliptic $K3$ surfaces. Michigan Math. J. {\bf 47} (2000), no. 3, 423--446.
%
%
\bibitem{Shimada} I. Shimada,
Rational double points on supersingular $K3$ surfaces.
Math. Comp. {\bf 73} (2004), no. 248, 1989--2017.


%
\bibitem{ShimadaOdd} I. Shimada,
Supersingular $K3$ surfaces in odd characteristic and sextic double planes. Math. Ann. {\bf 328} (2004), no. 3, 451--468.
%
\bibitem{ShimadaNormal} I. Shimada,
On normal $K3$ surfaces.
Michigan Math. J. {\bf 55} (2007), no. 2, 395--416. %
%
%\bibitem{ShimadaTransAMS} I. Shimada,
%Transcendental lattices and supersingular reduction lattices of a singular K3 surface.
%Trans. Amer. Math. Soc. {\bf 361} (2009), no. 2, 909?949.
%
\bibitem{MR3166075}
I. Shimada,
\newblock Projective models of the supersingular {$K3$} surface with {A}rtin
  invariant 1 in characteristic 5.
\newblock {J. Algebra}, \textbf{403} (2014), 273--299.

%
\bibitem{ShimadaZhang}
I. Shimada and De-Qi Zhang,
\newblock {$K3$} surfaces with ten cusps.
\newblock In {\em Algebraic geometry}, volume 422 of {\em Contemp. Math.},
  pages 187--211. Amer. Math. Soc., Providence, RI, 2007.
%
\bibitem{Shioda} T. Shioda,
Supersingular $K3$ surfaces.
Algebraic geometry (Proc. Summer Meeting, Univ. Copenhagen, Copenhagen, 1978), pp. 564--591,
Lecture Notes in Math., {\bf 732}.
%
%
\bibitem{Shioda2} T. Shioda,
On the Mordell-Weil lattices.
Comment. Math. Univ. St. Paul. {\bf 39} (1990), no. 2, 211--240.
%
\bibitem{MR0156890}
C.~T.~C. Wall,
\newblock Quadratic forms on finite groups, and related topics.
\newblock {Topology} \textbf{2} (1963), 281--298.

%
\bibitem{Yang1}
Jin-Gen Yang,
Sextic curves with simple singularities. Tohoku Math. J.
  (2) \textbf{48} (1996), no.~2, 203--227.
%
%
\bibitem{Yang2}
Jin-Gen Yang,
Enumeration of combinations of rational double points on quartic
  surfaces. Singularities and complex geometry (Beijing, 1994), AMS/IP Stud.
  Adv. Math., vol.~5, Amer. Math. Soc., Providence, RI, 1997, pp.~275--312.
%
%
\end{thebibliography}

\documentclass{amsart}

\usepackage{amssymb}
\usepackage{amsmath, longtable}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{_definition}[theorem]{Definition}
\newenvironment{definition}{\begin{_definition}\rm}{\end{_definition}}

\newtheorem{_remark}[theorem]{\it Remark}

\newtheorem{_claim}[theorem]{Claim}

\numberwithin{equation}{section}
\numberwithin{table}{section}
\numberwithin{figure}{section}

\renewcommand{\P}{\mathord{\mathbb  P}}
\newcommand{\Z}{\mathord{\mathbb Z}}

\newcommand{\Q}{\mathord{\mathbb Q}}

\newcommand{\TTT}{\mathord{\mathcal T}}

\newcommand{\set}[2]{\{\; {#1} \; \mid \; {#2} \;  \}}

\newcommand{\wt}{\widetilde}

\newcommand{\sprime}{\sp\prime}
\newcommand{\sptimes}{\sp\times}

\newcommand{\dual}{\sp{\vee}}

\newcommand{\sperp}{\sp\perp}

\newcommand{\OL}{\mathord{\mathcal{L}}}
\newcommand{\OLL}{\mathord{\tilde{\mathcal{L}}}}

\newcommand{\rmand}{\textrm{and}}

\newcommand{\quand}{\quad\rmand\quad}

\newcommand{\NS}{\mathord{\rm{NS}}}
\newcommand{\ch}{\operatorname{\mathord{\rm{char}}}\nolimits}
\newcommand{\disc}{\operatorname{\mathord{\rm{disc}}}\nolimits}
\newcommand{\Roots}{\mathord{\rm{roots}}}

\newcommand{\Hom}{\mathord{\rm{Hom}}}
\newcommand{\Aut}{\mathord{\rm{Aut}}}
\newcommand{\Sing}{\mathord{\rm{Sing}}}
\newcommand{\rank}{\operatorname{\mathord{\rm{rank}}}\nolimits}

\newcommand{\lat}{\Lambda}

\newcommand{\roots}{\Roots}
\newcommand{\Ex}{{\text{\rm Ex}}}

\begin{document}
\title[Supersingular $K3$ surfaces]{Dynkin diagrams of rank $20$ on supersingular $K3$ surfaces}

\author{Ichiro Shimada}
\address{
Department of Mathematics,
Graduate School of Science,
Hiroshima University,
1-3-1 Kagamiyama,
Higashi-Hiroshima,
739-8526 JAPAN
}
\email{shimada@math.sci.hiroshima-u.ac.jp
}

\author{De-Qi Zhang}
\address{
Department of Mathematics,
National University of Singapore,
10 Lower Kent Ridge Road,
Singapore 119076, SINGAPORE
}
\email{matzdq@nus.edu.sg
}
\begin{abstract}
We classify normal supersingular $K3$ surfaces $Y$ with total Milnor number $20$
in characteristic $p$,
where $p$ is an odd prime that does not divide the discriminant of  the
Dynkin type of the rational double points on $Y$.
This paper appeared in preprint form in the home page of the first author in the year 2005.
\end{abstract}

\subjclass{14J28}

\maketitle

\section{Introduction}
A \emph{Dynkin type} is, by definition,
a finite formal sum of the symbols $A_l (l\ge 1)$, $D_m (m\ge 4)$ and
$E_n (n=6,7,8)$ with non-negative integer coefficients.
For a Dynkin type $R$,
we denote by $L(R)$ the \emph{negative-definite} lattice
whose intersection matrix is $(-1)$ times the Cartan matrix of type $R$.
We denote by $\rank (R)$ the rank of $L(R)$, and by $\disc (R)$ the discriminant of $L(R)$.
\par
A \emph{normal $K3$ surface} is a normal surface whose minimal resolution is a $K3$ surface.
It is well-known that
a normal  $K3$ surface has only rational double points as its singularities
(\cite{Artin1962, Artin1966}).
Hence we can associate  a Dynkin type to the singular locus $\Sing (Y)$ of a normal $K3$ surface $Y$.
Recall that
the Milnor number of a rational double point of type $A_n$ (resp.~$D_n$,~resp.~$E_n$)
is $n$.
Hence
the rank of the Dynkin type of $\Sing (Y)$ is equal to
the sum of Milnor numbers of singular points on $Y$, that is, the \emph{total Milnor number of $Y$}.
In particular,
it is at most $21$.
\par
%
%
If the total Milnor number of a normal $K3$ surface $Y$ is $\ge 20$,
then the minimal resolution $X$ of $Y$ has Picard number $\ge 21$, and hence is a supersingular $K3$ surface
(in the sense of Shioda~\cite{Shioda}).
In~\cite[Theorem~3.7]{Goto}, Goto proved that a normal $K3$ surface $Y$ with total Milnor number
$21$ exists only when  the characteristic of the base field divides the discriminant of the Dynkin type of $\Sing (Y)$.
In~\cite{Shimada},
the first author made the complete list of the pairs $(R, p)$ of a Dynkin type $R$  of rank $21$ and a prime integer  $p$
such that
$R$ is  the  Dynkin type of the singular locus  of a normal $K3$ surface
in characteristic $p$.
\par
%
%
In this paper,
we investigate normal $K3$ surfaces with total Milnor number $20$.
\begin{definition}
Let $R$ be a Dynkin type of rank $20$.
A prime integer $p$ is called
an \emph{$R$-supersingular $K3$ prime}
if it satisfies the following:
\begin{itemize}
\item[(i)] $p$ is odd and does not divide $\disc (R)$, and
\item[(ii)] there exists a normal $K3$ surface $Y$ defined over an algebraically closed field
of characteristic $p$ such that $\Sing (Y)$ is of type $R$.
\end{itemize}
\end{definition}
The \emph{Artin invariant} of a supersingular $K3$ surface $X$ in characteristic $p$ is
the positive integer $\sigma$ such that the discriminant of the N\'eron-Severi lattice $\NS(X)$
of $X$ is equal to
$-p^{2\sigma}$. (See~\cite{Ar}).
We will prove that,
if $p$ is an $R$-supersingular $K3$ prime
for a Dynkin type $R$ with $\rank (R)=20$,
and if $Y$ is a normal $K3$ surface in the condition (ii) above,
then the Artin invariant of the minimal resolution of $Y$ is $1$.
It is known that,
for each $p$, the supersingular $K3$ surface with Artin invariant $1$ is unique up to isomorphisms
(\cite{Ogus, DKondo}).
Therefore the condition (i) and (ii) above is equivalent to (i) and the following:
\begin{itemize}
\item[(ii)${}\sprime$]
the supersingular $K3$ surface $X_p$ in characteristic $p$ with Artin invariant $1$
is birational to a normal $K3$ surface $Y$ such that  $\Sing (Y)$ is of type $R$.
\end{itemize}
In this paper, we present an algorithm to determine
the set of $R$-supersingular $K3$ primes for a given Dynkin type $R$ of rank $20$.
As a corollary, we prove the following.
\begin{theorem}\label{thm:Sigma1}
Let $R$ be a Dynkin type of rank $20$,
and let $a_R$ be the product of the odd prime divisors of $\disc (R)$.
We put $b_R:=8 a_R$ if $\disc (R)$ is even,
while $b_R:=a_R$ if $\disc (R)$ is odd.
Then there exists a subset $\Sigma_R$ of $(\Z/b_R\Z)\sptimes$
such that
a  prime integer $p$ is an $R$-supersingular $K3$ prime
if and only if
$p \bmod b_R\in \Sigma_R$.
\end{theorem}
In fact,
we have a result finer than above.
Let $Y$ be a normal supersingular $K3$ surface in characteristic $p\ne 2$
such that $\Sing (Y)$ is of Dynkin type $R$ with $\rank (R)=20$,
and let $X\to Y$ be the minimal resolution of $Y$.
We denote by $L_Y$ the sublattice of the N\'eron-Severi lattice $\NS (X)$ of $X$
generated by the classes of the exceptional curves of $X\to Y$.
Then $L_Y$ is isomorphic to $L(R)$.
Let $T_Y$ denote the orthogonal complement of $L_Y$ in $\NS (X)$.
Then $T_Y$ is an even indefinite lattice of rank $2$.
Our key observation is the following:
\begin{equation}\label{eq:observation}
tt\sprime\in p\Z\quad\textrm {for all}\quad t, t\sprime\in T_Y,
\end{equation}
where $tt\sprime\in \Z$ is the intersection number of the classes $t$ and $t\sprime$ in $\NS (X)$.
Thus we can define an indefinite lattice $T\sprime_Y$ of rank $2$ by
introducing a new bilinear form
$$
(t, t\sprime)_{T\sprime_Y}:=\frac{1}{p}(t t\sprime).
$$
on the $\Z$-module underlying $T_Y$.
It turns out that $\disc (T\sprime_Y)$ divides $\disc (R)$.
Note that, since $p$ is odd, $T\sprime_Y$ is an even lattice.
Let $\wt{L}_Y$ be the orthogonal complement of $T_Y$ in $\NS (X)$.
Then $\wt{L}_Y$ is an even overlattice of $L_Y$
such that the set $\Roots (\wt{L}_Y)$ of roots in $\wt{L}_Y$
coincides with the set $\Roots (L_Y)$ of roots in $L_Y$.
The following is a refinement of Theorem~\ref{thm:Sigma1}.
\begin{theorem}\label{thm:Sigma2}
Let $R$ be a Dynkin type of rank $20$,
let $T\sprime$ be an even indefinite lattice of rank $2$
such that $\disc (T\sprime)$ divides $\disc (R)$,
and let $\wt{L}$ be an even overlattice of $L(R)$ such that $\Roots (\wt{L})=\Roots (L(R))$.
Then there exist a subset $S_l$ of $\{1, -1\}$ for each odd prime divisor $l$ of $\disc (R)$,
and a subset $S_2$ of $\{1,3,5,7\}$,
such that  the following holds.
Let $p$ be an odd prime that does not divide $\disc(R)$.
Then there exists a normal $K3$ surface $Y$ in characteristic $p$
with $\Sing (Y)$ being of type $R$
such that $T\sprime_Y\cong T\sprime$ and $\wt{L}_Y\cong \wt{L}$
if and only if
\begin{equation}\label{eq:pcond}
\vtop{
\hbox{$\left(\frac{p}{l}\right)\in S_l$ for each odd prime divisor $l$ of $\disc(R)$, and}
\hbox{$p \bmod 8 \in S_2$.}
}
\end{equation}
If $\disc (R)$ is odd, then we have $S_2=\{1,3,5,7\}$.
\end{theorem}
Using computational algebra system {\tt Maple}, we have made
the complete list of $R$-supersingular $K3$ primes,
which  is
too large to be included in this paper.
It is available from the first author's home page
\begin{center}
\verb+http://www.math.sci.hiroshima-u.ac.jp/~shimada/K3.html+
\end{center}
in the plain text format.
From this list, we derive  the following fact:
\begin{theorem}\label{thm:half}
For each Dynkin type $R$ with $\rank (R)=20$,
the set of $R$-supersingular $K3$ primes is either empty or
has  a natural density $1/2$.
\end{theorem}
It would be very nice if Theorem~\ref{thm:half} is proved,
not by brute calculations of making the complete list, but by some
geometric reasonings.
\par
As another corollary of the key observation~\eqref{eq:observation},
we obtain the following:
\begin{corollary}\label{cor:degree}
Let $Y\subset \P^N$ be a normal supersingular $K3$ surface of total Milnor number $20$
such that $\Sing (Y)$ is of type $R$.
If the characteristic $p$ of the base field is odd and does not divide $\disc (R)$,
then the degree of $Y$ is divisible by $2p$.
\end{corollary}
Indeed the class of the pull-back of the hyperplane section of $Y$ to $X$ is
contained in $T_Y$.
Note that,
if $R$ is of rank $20$, then every prime divisor of $\disc(R)$ is $\le 19$.
Combining Corollary~\ref{cor:degree}
with~\cite[Theorem~3.7]{Goto},
we obtain the following:
\begin{corollary}\label{cor:mu19}
Let $Y$ be a normal $K3$ surface of degree $d$
in characteristic $p>19$.
If $p$ does not divide $d$,
then
the total Milnor number  of $Y$ is $\le 19$.
\end{corollary}
In particular, a sextic plane curve $C\subset \P^2$ or a quartic surface
$S\subset \P^3$  in characteristic $p>19$ with only rational double points as its singularities
has total Milnor number $\le 19$.
Yang~\cite{Yang1, Yang2} classified all possible configurations of rational double points
on sextic plane curves and quartic surfaces in characteristic $0$.
It would be interesting to investigate
Yang's  classification   in characteristic $p>19$.
See~\cite{ShimadaNormal} for a result on this problem.
\par
%
%
%
%
In our previous paper~\cite{ShimadaZhang},
we have proved that normal $K3$ surfaces with ten ordinary cusps
exist only in characteristic $3$.
This implies that
the set of $10A_2$-supersingular $K3$ primes is empty.
More generally, the proof of Dolgachev-Keum \cite[Lemma 3.2]{DK}
shows that, if $\disc (R)$ is a perfect square integer, then
there exist no $R$-supersingular $K3$ primes.
(See~Lemma~\ref{sqfree}.)
\par
There are $3058$ Dynkin types  of rank $20$.
Among them, there exist
$2437$ Dynkin types $R$ such that $\disc (R)$ is not a perfect square integer,
and $483$ Dynkin types with non-empty set of $R$-supersingular $K3$ primes.
\par
%
%
This paper is organized as follows.
In~Section~\ref{sec:lattice},
we reduce the problem of determining $R$-supersingular $K3$ primes
to the calculation of overlattices of $L(R)$ and their quadratic forms.
In~Section~\ref{sec:fqf},
we investigate how the multiplications by odd prime integers affects
the isomorphism classes of finite quadratic forms.
In~Section~\ref{sec:algorithm},
we  present an algorithm to calculate the set of $R$-supersingular $K3$ primes.
In the last section,
we explain the algorithm in detail by using an example.
\par
%
The study of the cases where $p$ is $2$ or divides $\disc (R)$
seems to need more subtle methods, and hence we do not treat these cases.
\par \vskip 1pc \noindent
{\bf Acknowledgement.}
We would like to thank the referees for the valuable suggestions and pointing out the reference
of C. T. C. Wall \cite{MR0156890} about quadratic forms.
\section{The N\'eron-Severi lattices of supersingular $K3$ surfaces}\label{sec:lattice}
A free $\Z$-module $\lat$ of finite rank with a non-degenerate
symmetric bilinear form
$\lat\times \lat \to \Z$
is called a \emph{lattice}.
Let $\lat$ be a lattice.
The \emph{dual lattice} $\lat\dual$ of $\lat$ is the $\Z$-module
$\Hom (\lat, \Z)$.
Then $\lat$ is naturally embedded into $\lat\dual$ as a submodule of finite index.
There exists a natural  $\Q$-valued symmetric bilinear form
on $\lat\dual$
that extends the $\Z$-valued symmetric bilinear form on $\lat$.
An \emph{overlattice} of $\lat$ is a submodule $N$ of $\lat\dual$
containing $\lat$
such that the bilinear form on $\lat\dual$
takes values in $\Z$ on $N\times N$.
If $\lat$ is a sublattice of a lattice $\lat\sprime$ with finite index,
then $\lat\sprime$ is embedded into $\lat\dual$ in a natural way,
and hence $\lat\sprime$ can be regarded as an overlattice of $\lat$.
\par
We say that $\lat$ is \emph{even} if $u^2\in 2\Z$ holds for every $u\in \lat$.
Let $\lat$ be an even negative-definite lattice.
A vector $r\in \lat$ is called a \emph{root}
if $r^2=-2$.
We denote by $\Roots (\lat)$ the set of roots in $\lat$.
Let $R$ be a Dynkin type.
Recall that $L(R)$ is the  negative definite root lattice of type $R$.
We put
\begin{eqnarray*}
\OLL(R)&:=&\set{\wt{L}}{\textrm{$\wt{L}$ is an even overlattice of $L(R)$}},\\
\OL(R)&:=&\set{\wt{L}\in \OLL(R)}{\Roots (\wt{L})=\Roots(L(R))}.
\end{eqnarray*}
Remark that we consider $\OLL(R)$ as a subset of
the set of submodules of $L(R)\dual$, and \emph{not} up to isometries of lattices.
\par
Let $D$ be a finite abelian group.
A \emph{quadratic form} $q$ on $D$ is, by definition,
a map $q: D\to \Q/2\Z$ that satisfies the following:
\begin{itemize}
\item[(i)] $q(nx)=n^2 q(x)$ for any $x\in D$ and any $n\in \Z$, and
\item[(ii)] the map $b[q]: D\times D\to \Q/\Z$ defined by
$$
b[q](x, y):= (q(x+y)-q(x)-q(y))/2
$$
is  a symmetric bilinear
form on $D$.
\end{itemize}
See Wall~\cite{MR0156890} for the complete classification of quadratic forms on finite abelian groups.
Let $q$ be a quadratic form on a finite abelian group $D$,
and let $H$ be a subgroup of $D$.
We put
$$
H\sperp :=\set{x\in D}{b[q](x, y)=0\;\;\textrm{for any}\; y\in H}.
$$
We say that $q$ is \emph{non-degenerate} if $D\sperp=\{0\}$.
\par
Let $\lat$ be an even lattice.
The finite abelian group $\lat\dual/\lat$ is called the \emph{discriminant group}
of $\lat$, and is denoted by $D_\lat$.
We define a quadratic form $q_\lat$ on $D_\lat$ by
$$
q_\lat (\bar x)=x^2\;\bmod\,2\Z\qquad (\textrm{where}\;\bar x:= x\bmod\,\lat\in D_\lat\;\textrm{for $x\in
\lat\dual$}),
$$
and call $q_\lat$ the \emph{discriminant form} of $\lat$.
It is easy to see that $q_\lat$ is non-degenerate, and that
$$
|D_\lat|=|\disc (\lat)|
$$
holds.
By Nikulin~\cite{Nikulin},
the map
$$
N\mapsto N/\lat
$$
induces a
 bijection  from the set of even overlattices of $\lat$ to
the set of isotropic subgroups of $(D_\lat, q_{\lat})$.
In particular,
we can calculate the set of even overlattices of
a given lattice by calculating the  set of isotropic subgroups
of its discriminant forms.
Let $N$ be an even overlattice of $\lat$.
Then we have a natural sequence of inclusions
$$
\lat\;\;\hookrightarrow\;\;N\;\;\hookrightarrow\;\; N\dual\;\; \hookrightarrow\;\; \lat\dual.
$$
Therefore $\disc (N)$ divides $\disc(\lat)$,
and the exponent of $D_N$ divides the exponent of $D_\lat$.
For a prime $p$,
we denote by $(D_\lat)_p$ and $(D_\lat)_{p\sprime}$ the $p$-part and the prime-to-$p$ part of
$D_\lat$, and by $(q_\lat)_p$ and $(q_\lat)_{p\sprime}$ the restrictions of $q$ to
$(D_\lat)_p$ and to $(D_\lat)_{p\sprime}$, respectively.
We have the following
orthogonal decomposition:
$$
(D_\lat, q_\lat)=((D_\lat)_p, (q_\lat)_p)\oplus ((D_\lat)_{p\sprime}, (q_\lat)_{p\sprime}).
$$
%
%
%
\par
We now state the main theorem of this section.
\begin{theorem}\label{Th2.1} 
Let $R$ be a Dynkin type with $\rank (R)=20$,
and let $p$ be an odd prime  that does not divide $\disc (R)$.
Then the following three conditions are equivalent:
\begin{itemize}
\item[(1)] $p$ is an $R$-supersingular $K3$ prime.
\item[(2)]
The unique supersingular $K3$ surface $X_p$ of Artin invariant
$1$
in characteristic $p$
is birational to a normal $K3$ surface $Y$ such that $\Sing(Y)$ is of Dynkin type $R$.
\item[(3)] There exist
an overlattice $\wt{L}\in \OL (R)$ and  a lattice $T\sprime$ of rank $2$ with signature $(1,1)$
such that $(D_{T\sprime}, p q_{T\sprime})$ is isomorphic to $(D_{\wt{L}}, -q_{\wt{L}})$,
where $p q_{T\sprime}$ is the discriminant form of $T\sprime$ multiplied by $p$.
\end{itemize}
\end{theorem}
%
%
%
%
%
%
A lattice $\Lambda$ is said to be \emph{$p$-elementary}
if its discriminant group $D_{\Lambda}$ is a $p$-elementary  group.
The following results due to
Artin~\cite{Ar} and Rudakov-Shafarevich~\cite{RS2}
reduce our geometric problem to calculations of lattices and finite quadratic forms.
\begin{theorem}[Artin~\cite{Ar} and Rudakov-Shafarevich~\cite{RS2}]
The N\'eron-Severi lattice of a supersingular $K3$ surface in characteristic $p$ is $p$-elementary.
\end{theorem}
Combining this result with the classification of indefinite lattices,
we have the following.
\begin{corollary}[Rudakov-Shafarevich~\cite{RS2}]
The isomorphism class of the N\'eron-Severi lattice $\NS(X)$ of a supersingular $K3$ surface $X$  is
uniquely
determined by the characteristic $p$ of the base field and the Artin invariant of $X$.
\end{corollary}
\begin{corollary}[Rudakov-Shafarevich~\cite{RS2}]\label{cor:RS2}
Suppose that $p$ is odd.
If $\Lambda$ is an even $p$-elementary
lattice of signature $(1, 21)$ and discriminant $-p^{2\sigma}$,
then $\Lambda$ is isomorphic to the N\'eron-Severi lattice of a supersingular $K3$ surface
in characteristic $p$ with Artin invariant $\sigma$.
\end{corollary}
%
%
%
The following easy result will be used in the proof of Theorem \ref{Th2.1}.
\begin{lemma}\label{TT'}
Let $T' = \Z t_1'\oplus \Z t_2'$ be a  lattice of rank $2$ with the intersection matrix
$(t_i'. t_j') = (t_{ij}')$.
Let $p$ be a prime that does not divide $\disc(T')$.
Define a lattice $T = \Z t_1\oplus  \Z t_2$ so that the intersection matrix
$(t_i . t_j) = (t_{ij})$ with $t_{ij} = p t_{ij}'$.
Then the following hold.
\begin{itemize}
\item
[(1)] $((D_T)_{p'}, (q_{T})_{p'}) \cong (D_{T\sprime}, p q_{T\sprime})$.
\item
[(2)] There exist positive integers $\ell_1$ and $\ell_2$ such that
$$T^{\vee}/T \cong \Z/(p \ell_1) \oplus \Z/(p \ell_2),
\hskip 0.5pc
(T')^{\vee}/T' \cong \Z/(\ell_1) \oplus \Z/(\ell_2).$$
\end{itemize}
\end{lemma}
\begin{proof}
Since $\rank(T') = 2$, we can write
$(T')^{\vee}/T' \cong \Z/(\ell_1) \oplus \Z/(\ell_2)$ so that $\ell_i > 0$, \
$\ell_1 | \ell_2$ and $\disc(T') = \det(t_{ij}') = \ell$
with $|\ell| = \ell_1 \ell_2$.
We can calculate the dual bases of $T'$ and $T$ as follows,
where $s_{11} = t_{22}'$, $s_{22} = t_{11}'$ and $s_{12} = s_{21} = -t_{12}' = -t_{21}'$:
$$((t_1')^*, (t_2')^*) = (t_1', t_2') (t_{ij}')^{-1} = \frac{1}{\ell} (t_1', t_2') (s_{ij}),$$
$$(t_1^*, t_2^*) = (t_1, t_2) (t_{ij})^{-1} = \frac{1}{p^2 \ell} (t_1, t_2) ( p s_{ij}).$$
Note that $(T^{\vee}/T)_{p'}$ is generated by
(cosets of) the two coordinates of the vector:
$$(p t_1^*, p t_2^*) = \frac{p}{\ell} (t_1, t_2) (s_{ij}).$$
Set $b_{T'} = b[q_{T'}]$, etc. Then
$$(b_{T'}((t_i')^* , (t_j')^*)) = (t_{ij}')^{-1} = \frac{1}{\ell}(s_{ij}),$$
$$(b_T(t_i^* , t_j^*)) = (t_{ij})^{-1} = \frac{1}{p^2 \ell}(p s_{ij}).$$
One can check that the following is an isomorphism of abelian groups:
$$(T')^{\vee}/T'  \rightarrow (T^{\vee}/T)_{p'}$$
$$ (t_i')^* + T' \mapsto p t_i^* + T.$$
Under the identification via this map, we have $p q_{T\sprime} = (q_T)_{p'}$.
This proves (1). Clearly, $\disc(T) = \det(t_{ij}) = p^2 \disc(T')$.
Also the expression of the dual basis shows that
$(T^{\vee}/T)_p$ is $p$-elementary. Thus (2) follows.
This proves the lemma.
\end{proof}
The following is the key to the proof of Theorem \ref{Th2.1}.
\begin{proposition}\label{Prop2.2}
Let $R$ be a Dynkin type with $\rank (R)=20$, and
let $L$ be an even overlattice of $L(R)$. Suppose that $p$ is an odd prime and that
$p \not| \ \disc(L)$.
\begin{itemize}
\item[(1)]
Suppose that $L \rightarrow \Lambda$ is a primitive embedding
into an even $p$-elementary lattice of signature $(1, 21)$
with non-cyclic $\Lambda^{\vee}/\Lambda$.
Let $T = L^{\perp}$ be the orthogonal complement
of $L$ in $\Lambda$.
Then {\rm (1a) $\sim$ (1e)} below hold.
\begin{itemize}
\item[(1a)]
$T$ is an even lattice of signature $(1, 1)$ such that $\disc(T) = -p^2 \disc(L)$ and
$T^{\vee}/T \cong \Z /(p \ell_1) \oplus \Z /(p \ell_2)$.
\item[(1b)]
There are a canonical isomorphism
$\varphi: L^{\vee}/L \rightarrow (T^{\vee}/T)_{p'}$
and the relation $(q_T)_{p'} = -q_L$ (after the
identification via $\varphi$).
\item[(1c)]
Write $T = \Z t_1 \oplus \Z  t_2$. Then $(t_i.t_j) = p(t_{ij}')$
for some $t_{ij}' \in \Z $.
\item[(1d)]
Let $T' = \Z t_1' \oplus \Z   t_2'$ be the lattice with the
intersection form $(t_i' . t_j') = (t_{ij}')$. Then
$(D_{T\sprime}, p q_{T\sprime}) \cong ((D_T)_{p'}, (q_{T})_{p'})
\cong (D_{L}, -q_{L})$.
\item[(1e)]
$\Lambda$ is the unique even $p$-elementary lattice of signature $(1, 21)$
and discriminant $-p^2$.
\end{itemize}
\item[(2)]
Conversely, suppose that $T$ is a lattice
of signature $(1, 1)$ satisfying
{\rm (1a) and (1b)}. Then there is a primitive embedding
$L \rightarrow \Lambda$ into the unique $p$-elementary even lattice $\Lambda$
of signature $(1, 21)$ and determinant $-p^2$
such that $T$  is isomorphic to the orthogonal complement $L^{\perp}$ of $L$ in $\Lambda$.
\end{itemize}
\end{proposition}
\begin{proof}
Consider the inclusions:
\begin{equation}\label{eq:second*}
 L \oplus T \subset \Lambda \subset \Lambda^{\vee}
\subset L^{\vee} \oplus T^{\vee}.
\end{equation}
Since $L \rightarrow \Lambda$ is a primitive embedding,
its dual is a surjection $\Lambda^{\vee} \rightarrow L^{\vee}$
which factors as $\Lambda^{\vee} \rightarrow L^{\vee} \oplus T^{\vee} \rightarrow L^{\vee}$
where the first is the inclusion in~\eqref{eq:second*} while the second is the projection.
This surjection induces surjections
$\Lambda^{\vee}/(L \oplus T) \rightarrow L^{\vee}/L$ and
$\varphi_1 : (\Lambda^{\vee}/(L \oplus T))_{p'} \rightarrow (L^{\vee}/L)_{p'} = L^{\vee}/L$
where the latter equality is because $\gcd(p, \disc(L)) = 1$.
Since $\Lambda$ is $p$-elementary we have $(\Lambda^{\vee}/\Lambda)_{p'} = 0$
and hence $(\Lambda^{\vee}/(L \oplus T))_{p'} = (\Lambda/(L \oplus T))_{p'}$.
Similarly, we have surjection
$\varphi_2 : (\Lambda^{\vee}/(L \oplus T))_{p'} \rightarrow (T^{\vee}/T)_{p'}$.
On the other hand, the inclusion~\eqref{eq:second*}  and the assumption that $\Lambda$ is $p$-elementary imply
$|L^{\vee}/L| \, |(T^{\vee}/T)_{p'}| = |(\Lambda/(L \oplus T))_{p'}|^2 \ge
|L^{\vee}/L| \, |(T^{\vee}/T)_{p'}|$ where the latter inequality is due to
the surjectivity of both $\varphi_i$. Thus both $\varphi_i$ are isomorphisms.
Set $\varphi = \varphi_2 \, \varphi_1^{-1}: L^{\vee}/L \rightarrow (T^{\vee}/T)_{p'}$.
For every $\overline{r'} \in L^{\vee}/L$, we write $\varphi(\overline{r'}) = \overline{t'}$,
and see that the coset of $r' + t'$
belongs to $(\Lambda/(L \oplus T))_{p'}$.
So $0 = q_{\Lambda}(\overline{r'+t'}) = q_R(\overline{r'}) + (q_{T})_{p'}(\overline{t'})$.
This proves (1b).
\par
Let $e$ be the exponent of the abelian group $L^{\vee}/L$ so that the latter is $e$-torsion.
This $e$ is coprime to $p$ by the assumption.
Then $\Lambda / (L \oplus T)$ ($\cong (L^{\vee} \oplus T^{\vee})/\Lambda^{\vee}$)
is $e$-torsion.
Indeed, for $r' + t' \in \Lambda \subset L^{\vee} \oplus T^{\vee}$,
we have, mod $L \oplus T$, that $e (r' + t') = e t' \in \Lambda \cap L^{\perp} = T$.
So $\Lambda/(L \oplus T)$ equals $(\Lambda/(L \oplus T))_{p'}$ and is isomorphic
to both $L^{\vee}/L$ and $(T^{\vee}/T)_{p'}$ via $\varphi_i$'s;
denote by $r$ the order of these three isomorphic groups,
which is coprime to $p$.
\par
We assert that $T^{\vee}/T$ is $pe$-torsion. Indeed, for
$t' \in T^{\vee}$, we have $e t' \in \Lambda^{\vee}$ and hence
$p e t' \in \Lambda \cap L^{\perp} = T$, because $\Lambda$ is $p$-elementary.
Since $T$ is of rank $2$, we have
$T^{\vee}/T \cong \Z /(p^{\varepsilon_1} \ell_1) \oplus \Z /(p^{\varepsilon_2} \ell_2)$
where $\ell_i \ge 1$,
each $\varepsilon_i \in \{0, 1\}$ and $\gcd(p, \ell_i) = 1$.
Note that $\Lambda^{\vee}/\Lambda \cong (\Z /(p))^{\oplus \lambda}$ for some
$\lambda \ge 2$.
The inclusion~\eqref{eq:second*} above implies that
$- p^{\varepsilon_1 + \varepsilon_2} \ell_1 \ell_2 \disc(L) =
\disc(T) \, \disc(L) = r^2 \disc(\Lambda) = -p^\lambda r^2$.
So $|\Lambda| = -p^2$ and both $\varepsilon_i = 1$.
Note also that $\ell_1 \ell_2 = |(T^{\vee}/T)_{p'}| = r = \disc(L)$.
This proves (1a) and (1e). The assertion (1c) is the observation
in (1a) that $(T^{\vee}/T)_p$ is $p$-elementary, $\rank(T) = 2$
and $\disc(T) = p^2 \times$ (some integer coprime to $p$)
and that the calculation of $T^{\vee}/T$ is essentially
the calculation of the matrix $(t_i . t_j)^{-1}$; see Lemma \ref{TT'}.
The assertion (1d) follows from (1b) and Lemma \ref{TT'}
by noting that $p$ does not divide $\disc(T') = -\disc(L)$.
\par
Next we prove Proposition \ref{Prop2.2} (2).
We define an overlattice $\Gamma$ of $L \oplus T$
by adding elements $r' + t' \in L^{\vee} \oplus T^{\vee}$ such that
$\varphi(\overline{r'}) = \overline{t'}$.
Note that $q_{L \oplus T}(\overline{r'+t'}) = q_L(\overline{r'}) + (q_T)_{p'}(\overline{t'}) = 0$,
so $\Gamma$ is an even overlattice of $L \oplus T$ such that the projections
induce isomorphisms: $L^{\vee}/L \cong \Gamma/(L \oplus T) \cong (T^{\vee}/T)_{p'}$.
Now $|\Gamma| = |L \oplus T|/\disc(L)^2 = \disc(T)/\disc(L) = -p^2$ by (1a).
Consider the inclusion~\eqref{eq:second*} above but with $\Lambda$ replaced by $\Gamma$,
we see that $\Gamma^{\vee}/\Gamma$ equals $(\Gamma^{\vee}/\Gamma)_p$, i.e., it is $p$-torsion
because so is $((L^{\vee} \oplus T^{\vee})/(L \oplus T))_p = (T^{\vee}/T)_p$ by (1a).
So $\Gamma$ is $p$-elementary. It is clear from the construction that
both $L \rightarrow \Gamma$ and $T \rightarrow \Gamma$ are primitive embeddings,
whence $T = L^{\perp}$ in $\Gamma$.
This proves Proposition \ref{Prop2.2}.
\end{proof}
\begin{proof}[Proof of Theorem \ref{Th2.1}]
We now prove Theorem \ref{Th2.1}, the direction $(1) \Longrightarrow (2)$.
So there is a normal $K3$ surface $Y$ defined over an algebraically closed
field $k$ with $\ch(k) = p$ such that $\Sing(Y)$ is of Dynkin
type $R$. Let $f : X \rightarrow Y$ be the minimal resolution and
$\Ex(f)$ the reduced exceptional divisor. Then $\Ex(f)$ is
also of Dynkin type $R$.
Since the Picard number $\rho(X) = \rho(Y) + \#\Ex(f) \ge 21$,
we have $\rho(X) = 22$ (see Artin \cite{Ar}) and hence $X$ is supersingular
in the sense of Shioda \cite{Shioda}. Let $\Lambda$ be the N\'eron-Severi lattice $\NS(X)$ of $X$.
Then
$\Lambda$ is $p$-elementary and $|\Lambda| = -p^{2 \sigma}$ where
$1 \le \sigma \le 10$ is the Artin invariant (Corollary~\ref{cor:RS2}).
Let $L$ denote
the sublattice of $\Lambda$ spanned by numerical equivalence classes of
irreducible components in $\Ex(f)$.
Then we have $L\cong L(R)$.
Let $\wt{L}$ be the closure of the sublattice $L$ in $\Lambda$.
Applying Proposition \ref{Prop2.2} to the primitive
embedding $\wt{L} \rightarrow \Lambda$, we see that
$\sigma = 1$. So Theorem \ref{Th2.1} (2) is true.
\par
Next we prove Theorem \ref{Th2.1}, the direction
$(2) \Longrightarrow (3)$. We use the notation above.
We assert that $\roots(\wt{L}) = \roots(L)$.
Indeed, suppose that $v \in \wt{L}$ is a $(-2)$-vector.
By considering $-v$ and the Riemann-Roch theorem, we may assume that
$v$ is represented by an effective divisor $V$ on $X_p$.
Since this $V$ is perpendicular to the pull back of an ample divisor
on $Y$, our $V$ is contractible to a point on $Y$, whence
$v = [V]$ belongs to $L$. The assertion is proved.
The rest of (3) follows from Proposition \ref{Prop2.2} applied
to the primitive embedding $\wt{L} \rightarrow \Lambda$.
\par
Finally, we prove Theorem \ref{Th2.1}, the direction
$(3) \Longrightarrow (1)$.
Define $T$ as in Lemma \ref{TT'}. Then
Propositions \ref{Prop2.2} (1a) and (1b) are satisfied
by $\wt{L}$ and $T$.
By Proposition \ref{Prop2.2} (both assertions there), there is a primitive embedding
$\wt{L} \rightarrow \Lambda$ into the unique even $p$-elementary
lattice of signature $(1, 21)$ and discriminant $-p^2$ such that $T = \wt{L}^{\perp}$
in $\Lambda$.
We have
$\NS(X_p) = \Lambda$.
Take a primitive element $v$ in $T^{\vee}$ such that $v^2 < -2$.
Let $h$ be a generator of $v^{\perp} \cap T^{\vee}$. So $h^2 > 0$.
We claim that
\begin{equation}\label{eq:rootss}
\roots(h^{\perp} \cap \Lambda) = \roots(\wt{L}).
\end{equation}
It is clear that the left-hand side of~\eqref{eq:rootss} includes the right-hand side of~\eqref{eq:rootss}.
Let $u$ be in the left-hand side of~\eqref{eq:rootss}. Write $u = r' + t'$ with $r' \in \wt{L}^{\vee}$
and $t' \in T^{\vee}$. Then $0 = h . u = h . t'$, whence
$t' \in T^{\vee} \cap h^{\perp} = \Z  [v]$.
So $t' = m v$ for some integer $m$. If $m \ne 0$, then
$-2 = u^2 = (r')^2 + (t')^2 \le m^2 v^2 < -2$, absurd.
So $m = 0$ and $u = r' \in \wt{L}^{\vee} \cap \Lambda = \wt{L}$
and hence $u \in$ the right-hand side of~\eqref{eq:rootss}. The claim is proved.
By considering $-h$ and isometry of $\Lambda$, we may assume
that a positive multiple of $h$ is represented by a nef and big Cartier
divisor $H$ on $X_p$
(see Rudakov-Shafarevich \cite{RS1}). Note that $|2H|$ is base point
free (see Nikulin \cite{Ni}, Proposition 0.1 and Saint-Donat \cite{SD},
Corollary 3.2). Let $f : X_p \rightarrow Y$ be the birational morphism
onto a normal surface, which is the Stein factorization of $\Phi_{|2H|}
: X_p \rightarrow \P^N$ with $N = \dim |2H|$.
Then $f$ is nothing but the contraction of all the curves
perpendicular to $H$. So by the genus formula and the Riemann-Roch theorem,
$\Ex(f)$ contains and consists of all curves representing elements in $\roots(h^{\perp} \cap \Lambda)
= \roots(\wt{L}) = \roots(L(R))$, whence $\Ex(f)$ is of Dynkin type $R$.
Thus $Y$ is a normal $K3$ surface with $\Sing(Y)$ of Dynkin type $R$.
Hence the assertion (1) is true. This completes the proof of Theorem \ref{Th2.1}.
\end{proof}
The following result
imposes on a Dynkin diagram $R$ of rank $20$
a necessary condition for the set of $R$-supersingular $K3$ primes to be non-empty.
The proof follows from the proof of Dolgachev-Keum \cite[Lemma 3.2]{DK}.
We reprove it here for the convenience of the readers.
\begin{lemma}\label{sqfree}
Let $R$  be a Dynkin diagram of rank $20$.
If the set of $R$-supersingular $K3$ primes is non-empty,
then $\disc(R)$ is not a perfect square.
\end{lemma}
\begin{proof}
By the assumption, there exist a $K3$ surface $X$
and $20$ smooth rational curves on $X$ whose numerical equivalence  classes span a
sublattice $L(R) \subset \NS(X)$ of Dynkin type $R$ and
which are contractible to rational double points on
a normal $K3$ surface $Y$. Since $\rho(X) \ge 1 + 20$,
we have $\rho(X) = 22$ and $X$ is supersingular.
Let $T$ denote the orthogonal complement of  $L(R)$ in $\NS(X)$.
Then $L(R) \oplus T$ is a sublattice of $\NS(X)$ of index $a$ say.
So we have:
\begin{equation}\label{eq:*}
\disc(R) \ \disc(T) = - a^2 p^{2 \sigma(X)},
\end{equation}
where $\disc(\NS(X)) = -p^{2 \sigma(X)}$ with
$\sigma(X) \in \{1, 2, \dots, 10\}$ the Artin invariant. 
\par
Suppose the contrary that $\disc(R)$ is a perfect square.
Then~\eqref{eq:*} implies that $- \disc(T)$ is a perfect square too.
By Conway-Sloane \cite{CS}, Chapter 15, Section 3, $T$ represents
zero: there is a non-zero vector $t$ in $T$ with $t^2 = 0$.
We may assume that $t$ is primitive in $T$.
By the Riemann-Roch theorem, we may assume that $t$ is represented
by an effective divisor. Applying Rudakov-Shafarevich \cite{RS2}, Chapter 3,
Proposition 3, there is a composite  $\sigma : \NS(X) \rightarrow \NS(X)$
of reflections with respect to $(-2)$-vectors
such that $\sigma(t)$ is represented by a general fibre $F$
of an elliptic or quasi-elliptic fibration $\varphi : X \rightarrow \P^1$.
There is a natural inclusion below where the lattice $F^{\perp}$ is the orthogonal in $\NS(X)$
of $\Z [F]$:
$$\sigma(L(R)) \rightarrow F^{\perp}/\Z [F].$$
Since $\rank(L(R)) = 20$, we can write
$F^{\perp}/\Z [F] \cong K_1 \oplus \dots \oplus K_r$
which includes $\sigma(L(R))$ as a sublattice of finite index $b$ say;
whence
\begin{equation}\label{eq:**}
 \disc(R) = b^2 \ \prod_{\ell=1}^r \disc(K_{\ell});
\end{equation}
moreover, each $K_{\ell}$ is of Dynkin type $A_{n(\ell)}$,
$D_{n(\ell)}$, or $E_{n(\ell)}$ so that $\varphi$
has reducible fibres of type $\widetilde{K_{\ell}}$ in the notation
of Cossec-Dolgachev \cite{CD};
see the reasoning below and the proof of Lemma 2.2 in Kondo~\cite{Ko}.
Let $j : J \rightarrow {\P}^1$ be the Jacobian fibration of $\varphi$
so that $j$ and $\varphi$ have the same type of singular fibres.
We note that $\rho(J) = 2 + 20$,
$J$ is supersingular, and $J$ has a torsion Mordell-Weil group $MW(j)$.
By Shioda \cite{Shioda2}, Theorem 1.3, we have
\begin{equation}\label{eq:***}
\prod_{\ell=1}^r \disc(K_{\ell}) = -p^{2 \sigma(J)} |MW(j)|^2,
\end{equation}
where $\disc(\NS(J)) = - p^{2 \sigma(J)}$ with $\sigma(J) \in \{1, 2, \dots, 10\}$.
Now~\eqref{eq:***} and~\eqref{eq:**} imply that $p$ divides $\disc(R)$, a contradiction.
Thus the lemma is proved.
\end{proof}
\section{Finite quadratic forms and prime integers}\label{sec:fqf}
Let $q$ and $q\sprime$ be  quadratic forms
on a finite abelian group $D$.
We denote by $d$ the order of $D$.
In this section,
we consider the set $K(q, q\sprime)$
of odd prime integers $p$ which are prime to $d$
such that $(D, pq)$ is isomorphic to $(D, q\sprime)$.
\par
%
%
For a prime $l$, we put
$$
T_l:=\begin{cases}
(\Z/8\Z)\sptimes  & \textrm{if $l=2$,}\\
\{1,-1\} & \textrm{if $l\ne 2$,}
\end{cases}
$$
and for an odd prime $p\ne l$, we define $\tau_l(p)\in T_l$ by
$$
\tau_l(p):=\begin{cases}
p\;\bmod 8 & \textrm{if $l=2$,}\\
\left(\frac{p}{l}\right) & \textrm{if $l\ne 2$.}
\end{cases}
$$
We then put
$$
T_d:=\prod_l T_l
$$
where $l$ runs through the prime divisors of $d$,
and put
$$
\tau_d(p):=(\tau_l(p))\;\;\in\;\; T_d
$$
for an odd  prime integer $p$ prime to $d$.
\begin{proposition}\label{prop:onlyfinite}
Let $p_1$ and $p_2$ be odd prime integers which are prime to $d$.
If $\tau_d(p_1)=\tau_d (p_2)$, then
saying $p_1\in K(q, q\sprime)$ is equivalent to saying $p_2\in K(q, q\sprime)$
\end{proposition}
\begin{proof}
It is enough to prove that $(D, p_1q)$ and  $(D, p_2 q)$ are isomorphic.
Let $l$ be an odd prime divisor of $d$,
and let $\nu$ be the largest  integer such that $l^\nu | d$.
It follows from $\tau_l (p_1)=\tau_l(p_2)$
that there exists an integer $a_l$
such that $p_1\equiv a_l^2 p_2\bmod\,l^\nu$ holds.
Note that $a_l$ is prime to $l$.
Suppose that $d$ is even.
It follows from $\tau_2 (p_1)=\tau_2(p_2)$
that there exists an integer $a_2$
that satisfies $p_1\equiv a_2^2 p_2\bmod\,2^{\mu+1}$,
where $\mu$ is the largest integer such that $2^\mu|d$.
Note that $a_2$ is odd.
By the Chinese Remainder Theorem,
we have an integer $a$  that satisfies
$a\equiv a_l\bmod\, l^\nu$ for each odd prime divisor $l$  of $d$,
and
$$
a\equiv\begin{cases}
a_2\bmod\, 2^{\mu+1} &\textrm{if $d$ is even,}\\
\rlap{1}\phantom{a_2}\bmod\, 2 &\textrm{if $d$ is odd.}
\end{cases}
$$
Then we have
 $p_1\equiv a^2 p_2 \bmod\, 2d$.
 Note that $a$ is prime to $d$.
 Since $b[q](x, x)=q(x) \bmod\,\Z$,
 $q(x)$ is contained in $(1/d)\Z/2\Z\subset\Q/2\Z$
 for any $x\in D$.
 Therefore we have
 $$
 p_1 q=a^2 p_2 q.
 $$
 The multiplication by $a$ induces an automorphism $\alpha $ of $D$.
 Since $\alpha^* ( p_2 q)=p_1 q$,
 we see that $p_1q$ and $p_2 q$ are isomorphic.
\end{proof}
\begin{corollary}
There exists a subset $S(q, q\sprime)$ of $T_d$ such that
$K(q, q\sprime)$ is equal to the set of odd prime integers $p$ which are prime to $d$
such that $\tau_d (p)\in S(q, q\sprime)$.
\end{corollary}
%
%
%
%
%
%
%
%
%
%
%
\section{Algorithm}\label{sec:algorithm}
Let $R$ be a Dynkin type of rank $20$.
We put
$$
T(R):=\prod_{l} T_l,
$$
where $l$ runs through the set of prime divisors of $\disc (R)$,
and for an odd prime $p$ that does not divide $\disc (R)$,
we put
$$
\tau (p):= (\tau_l (p))_l\in T(R).
$$
In this section, we present an algorithm to obtain a subset
$S(R)\subset T(R)$ with the following property:
an odd prime $p$ that does not divide $\disc (R)$ is an $R$-supersingular $K3$ prime
if and only if $\tau (p)\in S(R)$.
\subsection{Step 1}
We first calculate the  set $\OLL(R)$ of even overlattices of $L(R)$ using
Proposition 1.4.1 of  Nikulin~\cite{Nikulin}.
For each even overlattice $\wt{L}$ of $L(R)$, we can calculate the set $\Roots (\wt{L})$  of roots
of $\wt{L}$ by the method described in~\cite{ShimadaEll},~\cite{ShimadaOdd} or~\cite{MR3166075}.
Comparing $\Roots (\wt{L})$
with $\Roots (L(R))$ for each $\wt{L}$,
we make the set $\OL (R)$.
\subsection{Step 2}
We calculate the discriminant group $D_{\wt{L}}$
for each $\wt{L}\in \OL (R)$,
and make the set $\OL\sprime (R)$ of all $\wt{L}\in \OL (R)$
such that the length of $D_{\wt{L}}$
 is $\le 2$.
For each $\wt{L}\in \OL\sprime (R)$,
we calculate the isomorphism class of the
finite quadratic form $(D_{\wt{L}}, -q_{\wt{L}})$.
\subsection{Step 3}
For each $\wt{L}\in \OL\sprime  (R)$, we do the following calculation.
We put $d:=\disc (\wt{L})$,
which is a positive integer.
First we make the list $\TTT(d)$ of isomorphism classes of even indefinite lattices
 $T\sprime$ of rank $2$
with discriminant $-d$
using the classical theory of binary forms due to  Gauss.
(See~\cite[Chapter 15, \S3.3]{CS}.)
For each $T\sprime\in \TTT(d)$ we calculate the discriminant group $D_{T\sprime}$ of $T\sprime$.
If $D_{T\sprime}$ is isomorphic to $D_{\wt{L}}$,
then we calculate the set
$$
S(\wt{L}, T\sprime):=\prod_{l|\disc (R)} S_l(\wt{L}, T\sprime)\;\;\subset\;\; T(R)
$$
such that $(D_{T\sprime}, pq_{T\sprime})$ is isomorphic to $(D_{\wt{L}}, -q_{\wt{L}})$
if and only if $\tau _l (p)\in S_l(\wt{L}, T\sprime)$ for each
prime divisor $l$ of $\disc (R)$.
In virtue of Proposition~\ref{prop:onlyfinite},
we have to check only a finite number of prime integers.
(Note that
the set of prime divisors of $|D_{\wt{L}}|$ is a subset of the set of prime divisors of $\disc (R)=|D_{L(R)}|$.
If a prime divisor $l$ of $\disc (R)$
does not divide  $\disc (\wt{L})$, then we put $S_l(\wt{L}, T\sprime)=T_l$.)
If $D_{T\sprime}$ is not isomorphic to $D_{\wt{L}}$,
then we put $S(\wt{L}, T\sprime)=\emptyset$.
\par
%
%
The set $S(R)$ is the union of all $S(\wt{L}, T\sprime)$,
where $\wt{L}$ runs through the set $\OL\sprime (R)$ and $T\sprime$ runs through the set
$\TTT(\disc (\wt{L}))$.
\section{Example}\label{sec:detail}
\newcommand{\diag}{\mathord{\rm diag}}
We will explain the case
$$
R:=D_7+A_{11}+2 A_1
$$
in detail.
The discriminant form of
the negative-definite root lattice $L(R)$ is expressed by the diagonal matrix
$$
\diag[-7/4, -11/12, -1/2, -1/2]
$$
with respect to  the basis of the discriminant group
$$
D_{L(R)}\;\;\cong\;\; \Z/4\Z\;\times\; \Z/12\Z\;\times\; \Z/2\Z\;\times\;\Z/2\Z
$$
given    in~\cite[Section 6]{ShimadaEll}.
There are eight  isotropic vectors in $D_{L(R)}$:
$$
\begin{array}{l}
0:=[0,0,0,0],\;\;
v_1:=[0,6,1,1],\;\;
v_2:=[1,3,0,0],\;\;
v_3:=[1,9,0,0],\;\;
v_4:=[2,0,1,1],\\
v_5:=[2,6,0,0]=2v_2=2v_3,\;\;
v_6:=[3,3,0,0]=-v_3,\;\;
v_7:=[3,9,0,0]=-v_2.
\end{array}
$$
Let $L_{(i)}$ be the even overlattice of $L(R)$ corresponding to
the totally isotropic subgroup
of $D_{L(R)}$ generated by $v_i$.
The Dynkin type of $\Roots (L_{(i)})$
is equal to $R$ if $i\ne 4$,
while it is $A_{11}+D_9$ if $i=4$.
Hence the even overlattice $L(H)$ of $L(R)$
corresponding to an totally isotropic subgroup
$H$ satisfies $\Roots (L(H))=\Roots(L(R))$ if and only if
$v_4\notin H$.
The totally isotropic subgroups that do not contain $v_4$ are listed as follows:
$$
H_0=\{0\},\;\;
H_1=\{0, v_1\},\;\;
H_2=\{0, v_5\}, \;\;
H_3=\{0, v_2, v_5, v_7\}, \;\;
H_4=\{0, v_3, v_5, v_6\}.
$$
Let $\gamma\in \Aut(L(R))$ be the isometry of $L(R)=L(D_7)\oplus L(A_{11}+2A_{1})$
that is the multiplication by $-1$ on the factor $L(D_7)$
and the identity on $L(A_{11}+2A_{1})$.
Then the action of $\gamma$ on $D_{L(R)}$ interchanges $H_3$ and $H_4$.
Therefore the even lattices $L(H_3)$ and $L(H_4)$ are isomorphic.
The lengths of $D_{L(H_0)}$ and $D_{L(H_2)}$ are $\ge 3$,
while the lengths of $D_{L(H_1)}$ and $D_{L(H_3)}\cong D_{L(H_4)}$ are $2$.
\par
%
%
The discriminant  form of $L(H_1)$
multiplied by $(-1)$ is given by
$$
\left(\Z/4\Z\times \Z/4\Z,
\left[\begin{array}{cc}7/4 & 0 \\ 0 & 3/4 \end{array}\right]\right)
\times
\left(\Z/3\Z,
\left[\begin{array}{c}
2/3
\end{array}
\right]
\right).
$$
There are four isomorphism classes of even indefinite lattices
of rank $2$ with discriminant $-48$:
$$
S_\pm:=\left[
\begin{array}{cc}
\pm 2 & 6 \\
6 & \mp 6
\end{array}
\right],
\qquad
T_\pm:=\left[
\begin{array}{cc}
\pm 4 & 4 \\
4 & \mp 8
\end{array}
\right].
$$
The discriminant group of $S_\pm$ is isomorphic to $\Z/2\Z\times \Z/8\Z\times\Z/3\Z$,
and hence is not isomorphic to $D_{L(H_1)}$.
The discriminant forms  of $T_\pm$ are
$$
\left(\Z/4\Z\times \Z/4\Z,
\left[\begin{array}{cc}\pm 1/4 & 0 \\ 0 & \pm 5/4 \end{array}\right]\right)
\times
\left(\Z/3\Z,
\left[\begin{array}{c}
\pm 2/3
\end{array}
\right]
\right).
$$
Hence we have the following equivalence for prime integers $p\ne 2, 3$:
$$
\begin{array}{rcl}
p\in K(-q_{L(H_1)}, q_{T_+})
&\Longleftrightarrow
&p\bmod 8\equiv  3\;\textrm{or}\;7\quand(\frac{p}{3})=1,\\
p\in K(-q_{L(H_1)}, q_{T_-})
&\Longleftrightarrow
&p\bmod 8\equiv  1\;\textrm{or}\;5\quand(\frac{p}{3})=-1.
\end{array}
$$
There are two isomorphism classes of even indefinite lattices
of rank $2$ with discriminant $-12$:
$$
U_\pm:=\left[
\begin{array}{cc}
\pm 2 & 2 \\
2 & \mp 4
\end{array}
\right].
$$
By the same calculation as above,
 we have the following equivalence for prime integers $p\ne 2, 3$:
$$
\begin{array}{rcl}
p\in K(-q_{L(H_3)}, q_{U_+})
&\Longleftrightarrow
&p\bmod 8\equiv  1\;\textrm{or}\;5\quand(\frac{p}{3})=-1,\\
p\in K(-q_{L(H_3)}, q_{U_-})
&\Longleftrightarrow
&p\bmod 8\equiv  3\;\textrm{or}\;7\quand(\frac{p}{3})=1.
\end{array}
$$
Thanks to the equalities
$$
K(-q_{L(H_1)}, q_{T_+})= K(-q_{L(H_3)}, q_{U_-}),
\quad
K(-q_{L(H_1)}, q_{T_-})=K(-q_{L(H_3)}, q_{U_+}),
$$
the natural density of $R$-supersingular $K3$ primes is $1/2$.
\end{document}